At 2:00 p.m. two cars start toward each other from towns 240 miles apart. If the rate of one car is 10 mph faster than the other, how fast does each car go,of they meet at 5:00p.m.?
Answer by checkley75(3666)

THEY MEET AFTER THREE HOURS FROM THE START & EACH TRAVEL FOR 3 HOURS. ONE TRAXELS @ X MPH & THE OTHER @ (X+10) MPH.
240=3X+3(X+10)
240=3X+3X+30
6X=240-30
6X=210
X=210/6
X=35 MPH FOR THE SLOWER CAR.
35+10=45 MPH FOR THE FASTER CAR.
PROOF
240=3*35+3*45
240=105+135
240=240